• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Higinbotham

    Log entry time 15:57:46 on November17,2005

    Entry number 160461

    keyword=Shift Summary


    Shift Summary
    -------------

    The first half of the shift was spent waiting for MCC to get the
    beam back to the state it was in before last night's storm. Around 12:30 good beam was restored. Since then we have been taking good data.

    Happex Run Number
    -----------------

    Time Run Number Pairs Note
    0832:0935 6509 12.8k Many Trips
    0950:1006 6510 ?0.0k Many Trips
    1154:1209 6511 1.3k Compton Test
    1230:1335 6512 31.5k Stable Beam :-)
    1335:1430 6513 31.6k No Dither Cycle w/o Trip
    1430:1525 6514 40.1k
    15:30 thru end of shift

    Compton Run Number
    ------------------

    Time Run Number Polarization [%]
    1230:1430 12897 89 +/- 1
    1435: 12898

    Raster Check
    ------------

    Time Run Number
    1340 3089